OpenStudy (anonymous):

Lin Alg question... L: R^3 --> R^2 defined by L((x,y,z)) =xz ..is it LINEAR or NOT LINEAR..explain the steps i'm pretty stuck on it! Thx. D55.

OpenStudy (anonymous):

really all i know is L(x +y +z)= Lx + Ly + Lz and L(cx) = cL(u) if its linear... dont really know what to do with that tho

OpenStudy (anonymous):

L(cx)=cL(x)**

OpenStudy (anonymous):

\(L((x_1,y_1,z_1)+(x_2,y_2,z_2))=L((x_1+x_2,y_1+y_2,z_1+z_2))=(x_1+x_2)(z_1+z_2)\). \(L(x_1,y_1,z+1)+L(x_2,y_2,z_2)=x_1z_1+x_2z_2\ne(x_1+x_2)(z_1+z_2)\) . So it's NOT lineaer.

OpenStudy (anonymous):

There is a typo in the second line. It should be \(z_1\) not z+1.

OpenStudy (anonymous):

thanks anwar appreciate it once again!

OpenStudy (anonymous):

You're welcome :)

OpenStudy (anonymous):

donny r u here?